How we can find the length of cirrumradius and inner radius of a triangle. 

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
Circum radius = R = a/2 sinA = b/ 2 sinB = c/2 sinC
 
Inner radius = r = Area of traingle /(semi perimeter)